# Lattice Boltzmann methods for soft flowing matter

### v1

@@ VI.2.3 Learning near-contact pseudo-potentials
In Ref.Montessori et al. (2019a) an effective way to incorporate disjoining, near-contact interactions within a chromodynamic LB model for multicomponent flows is presented, which stands as an extreme instance of coarse-graining of the complex physics arising when two fluid interfaces covered by a surfactant come into close contact. Such a picture is typical soft many-body flowing systems, such as dry foams and dense emulsions. As mentioned in Section II, such near-contact interactions aim at condensing a plethora of effects due to the onset of forces arising at nanometric scales, such as electrostatic double-layer, dispersion forces and Casimir-like interactions Derjaguin and Landau (1941); Verwey and Overbeek (1948), within a simplified, one-parameter density functional.
@@
The hard way to solve such issues is to systematically ”down-scale” the problem, by studying the details of the interactions through micro-scale techniques, such as molecular dynamics and perhaps even electronic structure methods in case electronic degrees of freedom take an active role Bocquet and Charlaix (2010); Coquinot et al. (2023). The main problem of such a microscopic approach is the lack of computer power to reach out to scales of experimental and engineering design interest, say centimeters. The alternative, top-down approach, is to augment mesoscale forces and pseudo-potentials with additional terms which would retain the essential microphysics within a viable computational framework. Translated into formulas, the near-contact force per unit volume along direction $x_{a}$ is given by:
$F_{a,nci}=\nabla\cdot\Sigma_{nci}=\nabla\cdot\left(A_{nci}[z;\kappa_{int},u_{rel}]\frac{\nabla\phi}{|\nabla\phi|}\right)\vec{n}\delta_{ab}$ (93)
where $z$ is the spatial coordinate along the normal $\vec{n}$ of a given interfaces, $u_{rel}$ is the relative velocity and $\kappa_{int}$ is the local curvature of the interface.
@@ VI.3 Machine learning for the deformable many-body problem
Figure 25: (a) A picture of a sequence of monodisperse droplets in a divergent microfluidic channel. This represents an example of a training dataset for YOLO. Each droplet is identified by a bounding box, whose dimensions are noted in an associated file. (b) Examples of DropTrack’s outputs, which identify and track droplets of a dense emulsion flowing in microchannels of different geometry. Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 license. No changes were made to the original figures in Ref. Durve et al. (2023).
@@ VI.3.1 Learning shape dynamics
In many soft matter applications, there is substantial scope for controlling not only the position and velocity of the single droplets, but also their shape changes as such droplets move along the flow. Since this adds a significant number of internal degrees of freedom, machine learning can find profitable use to accomplish this task as well. For instance, Li et al. Li et al. (2024) used machine learning techniques to predict the shape of bubble droplets to be used in a multiphase LB simulations using the pseudopotential method. A regularized feedforward deep neural network, made of a first hidden layer of fifty neurons and a second one of ten neurons, was then trained on this data with a gradient descent method, providing a R-coefficient ranging between 0.9965 and 1. However, the study is limited by the narrow range of surface tension (0.01 and 0.02) and viscosity ratios (0.05 and 0.10), as well as by the bi-dimensionality of the simulation. Recently, a class of machine learning methodologies, called autoencoder, has been employed to acquire minimal droplet shape descriptors, subsequently applied in forecasting droplet breakup occurrences Khor et al. (2019); Durve et al. (2024b), a process crucial in many industrial applications Rosenfeld et al. (2014b); Gai et al. (2016).
@@ VI.4 Prospects for quantum computing of soft fluids
We finally discuss a further perspective about potential links between lattice Boltzmann and quantum computing. Indeed, quantum computing is one of the most vibrant topics of modern science, holding promises of spectacular applications far beyond the reach of classical electronic computers, if only for a limited set of applications Arute et al. (2019); Ippoliti et al. (2021). The main point is that qubits, the quantum analog of classical bits, represent an arbitrary superposition of the two fundamental states $|0>$ (ground state) and $|1>$ (excited state), so that a collection of $Q$ qubits spans an exponentially large Hilbert space, consisting of $2^{Q}$ classical states. The blue-sky scenario for fluids is mind-boggling. Given that a fluid flow at Reynolds number $Re$ consists of about $Re^{3}$ active degrees of freedom, the number of qubits required to represent such state-space is given by
@@
The idea is very simple and best illustrated by means of a zero-dimensional nonlinear system, namely the logistic equation:
$\frac{dx}{dt}=x(1-Rx),\;\;\;x(t=0)=x_{0}$ (101)
@@
Figure 26: Left panel: Initial conditions, source and solutions of the velocity field of the Burger equation computed at Re=20 with a spatial discretization of 16 grid points and run on a classical computer. Right panel: $l_{2}$ norm of the absolute error between different Carleman truncation levels. Note that, the error of the Carleman method converges exponentially with $N$. Copyright (2021) National Academy of Sciences Ref.Liu et al. (2021).
@@
It should be mentioned that much better convergence can be obtained by applying the Carleman procedure to the Lattice Boltzmann equation Itani and Succi (2022); Succi et al. (2023); Sanavio and Succi (2023). Unfortunately, the corresponding Carleman-LB algorithm is non-local, which reflects into an unviable depth of the corresponding quantum circuit. As a result, with the current state of affairs, the implementation of a quantum algorithm for fluid flows is still open. Coming back to soft flowing matter, the good news is that the Reynolds number is generally moderate, often below $100$, hence, once the aforementioned problems are solved (assuming they will), about a hundred effective qubits will suffice to quantum simulate complex states of soft flowing matter. Work along this line is currently in progress.
@@ VII Conclusions
Summarizing we have discussed a series of Lattice Boltzmann methods to simulate microscale multi-component soft flows under strong geometric confinement. Such methods leverage the flexibility of the LB method to efficiently incorporate mesoscale physics beyond hydrodynamics and yet in no need of detailed molecular specificity. The inclusion of near-contact interactions via a simple repulsive force stands as a paradigmatic example of this approach and has proven successful for a broad variety of complex rheological applications which would be very hard to treat with different methods, particularly in the case of dense confined emulsions.
@@ Acknowledgements
A.T., M.D., M.L., A.M., and S.S. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 Framework Program (No. FP/2014-2020) ERC Grant Agreement No. 739964 (COPMAT) and ERC-PoC2 grant No. 101081171 (DropTrack). A.T. and M.L. acknowledge the support of the Italian National Group for Mathematical Physics of INdAM (GNFM-INdAM). J.M.T thanks the FRQNT ”Fonds de recherche du Québec - Nature et technologies (FRQNT)” for financial support (Research Scholarship No. 314328).

### v2

@@ VI.2.3 Learning near-contact pseudo-potentials
In Ref.Montessori et al. (2019a) an effective way to incorporate near-contact interactions within a chromodynamic LB model for multicomponent flows is presented, which stands as an extreme instance of coarse-graining of the complex physics arising when two fluid interfaces covered by a surfactant come into close contact. Such a picture is typical of soft many-body flowing systems, such as dry foams and dense emulsions. As mentioned in Section II, such near-contact interactions aim at condensing a plethora of effects due to the onset of forces arising at nanometric scales, such as electrostatic double-layer, dispersion forces and Casimir-like interactions Derjaguin and Landau (1941); Verwey and Overbeek (1948), within a simplified, one-parameter density functional.
@@
The hard way to solve such issues is to systematically ”down-scale” the problem, by studying the details of the interactions through micro-scale techniques, such as molecular dynamics and perhaps even electronic structure methods in case electronic degrees of freedom take an active role Bocquet and Charlaix (2010); Coquinot et al. (2023). The main problem of such a microscopic approach is the lack of computer power to reach out to scales of experimental and engineering design interest, say centimeters. The alternative, top-down approach, is to augment mesoscale forces and pseudo-potentials with additional terms which would retain the essential microphysics within a viable computational framework. Translated into formulas, the near-contact force per unit volume along direction $x_{a}$ is given by
$F_{a,nci}=\nabla\cdot\Sigma_{nci}=\nabla\cdot\left(A_{nci}[z;\kappa_{int},u_{rel}]\frac{\nabla\phi}{|\nabla\phi|}\right)\vec{n}\delta_{ab},$ (93)
where $z$ is the spatial coordinate along the normal $\vec{n}$ of a given interface, $u_{rel}$ is the relative velocity and $\kappa_{int}$ is the local curvature of the interface.
@@ VI.3 Machine learning for the deformable many-body problem
Figure 27: (a) A picture of a sequence of monodisperse droplets in a divergent microfluidic channel. This represents an example of a training dataset for YOLO. Each droplet is identified by a bounding box, whose dimensions are noted in an associated file. (b) Examples of DropTrack’s outputs, which identify and track droplets of a dense emulsion flowing in microchannels of different geometry. Copyright granted under the CC 4.0 license. No changes were made to the original figures in Ref. Durve et al. (2023).
@@ VI.3.1 Learning shape dynamics
In many soft matter applications there is substantial scope for controlling not only the position and velocity of the single droplets, but also their shape changes as such droplets move along the flow. Since this adds a significant number of internal degrees of freedom, machine learning can find profitable use to accomplish this task as well. For instance, Li et al. Li et al. (2024) used machine learning techniques to predict the shape of bubble droplets to be used in a multiphase LB simulations using the pseudopotential method. A regularized feedforward deep neural network, made of a first hidden layer of fifty neurons and a second one of ten neurons, was then trained on this data with a gradient descent method, providing a R-coefficient ranging between 0.9965 and 1. However, the study is limited by the narrow range of surface tension (0.01 and 0.02) and viscosity ratios (0.05 and 0.10), as well as by the bi-dimensionality of the simulation. Recently, a class of machine learning methodologies, called autoencoder, has been employed to acquire minimal droplet shape descriptors, subsequently applied in forecasting droplet breakup occurrences Khor et al. (2019); Durve et al. (2024b), a process crucial in many industrial applications Rosenfeld et al. (2014b); Gai et al. (2016).
@@ VI.4 Prospects for quantum computing of soft fluids
We finally discuss a further perspective about potential links between LB methods and quantum computing. Indeed, quantum computing is one of the most vibrant topics of modern science, holding promises of spectacular applications far beyond the reach of classical electronic computers, if only for a limited set of applications Arute et al. (2019); Ippoliti et al. (2021). The main point is that qubits, the quantum analog of classical bits, represent an arbitrary superposition of the two fundamental states $|0>$ (ground state) and $|1>$ (excited state), so that a collection of $Q$ qubits spans an exponentially large Hilbert space, consisting of $2^{Q}$ classical states. The blue-sky scenario for fluids is mind-boggling. Given that a fluid flow at Reynolds number $Re$ consists of about $Re^{3}$ active degrees of freedom, the number of qubits required to represent such state-space is given by
@@
The idea is very simple and best illustrated by means of a zero-dimensional nonlinear system, namely the logistic equation
$\frac{dx}{dt}=x(1-Rx),\;\;\;x(t=0)=x_{0},$ (101)
@@
Figure 28: Left panel: Initial conditions, source and solutions of the velocity field of the Burger equation computed at Re=20 with a spatial discretization of 16 grid points and run on a classical computer. Right panel: $l_{2}$ norm of the absolute error between different Carleman truncation levels. Note that the error of the Carleman method converges exponentially with $N$. Copyright (2021) National Academy of Sciences Ref.Liu et al. (2021).
@@
It should be mentioned that much better convergence can be obtained by applying the Carleman procedure to the LB equation Itani and Succi (2022); Succi et al. (2023); Sanavio and Succi (2023). Unfortunately, the corresponding Carleman-LB algorithm is non-local, which reflects into an unviable depth of the corresponding quantum circuit. As a result, with the current state of affairs, the implementation of a quantum algorithm for fluid flows is still open. Coming back to soft flowing matter, the good news is that the Reynolds number is generally moderate, often below $100$, hence, once the aforementioned problems are solved (assuming they will), about a hundred effective qubits will suffice to quantum simulate complex states of soft flowing matter. Work along this line is currently in progress.
@@ VII Conclusions
Summarizing we have discussed a series of LB methods to simulate microscale multi-component soft flows under strong geometric confinement. Such methods leverage the flexibility of the LB method to efficiently incorporate mesoscale physics beyond hydrodynamics and yet in no need of detailed molecular specificity. The inclusion of near-contact interactions via a simple repulsive force stands as a paradigmatic example of this approach and has proven successful for a broad variety of complex rheological applications which would be very hard to treat with different methods, particularly in the case of dense confined emulsions.
@@ Acknowledgements
A.T., M.D., M.L., A.M., and S.S. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 Framework Program (No. FP/2014-2020) ERC Grant Agreement No. 739964 (COPMAT) and ERC-PoC2 grant No. 101081171 (DropTrack). A.T. and M.L. acknowledge the support of the Italian National Group for Mathematical Physics of INdAM (GNFM-INdAM). J.M.T thanks the FRQNT “Fonds de recherche du Québec - Nature et technologies (FRQNT)” for financial support (Research Scholarship No. 314328).